OpenStudy (anonymous):

Expand the logarithmic expression. log7(n/5)

Nnesha (nnesha):

quotient rule\[\large\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

\[\huge\rm log_7 \frac{ n }{ 5}\]which property you should apply ?

OpenStudy (anonymous):

quotient rule

OpenStudy (usukidoll):

yes!

Nnesha (nnesha):

yep right btw 7 is base right ?

OpenStudy (anonymous):

yes 7 is the base

Nnesha (nnesha):

alright

Nnesha (nnesha):

so how would you expand that ?

OpenStudy (anonymous):

log7(5)-log5(n)

Nnesha (nnesha):

hmm log5(n) ??

OpenStudy (anonymous):

oh, i meant 7

OpenStudy (usukidoll):

we need to have the same base value... 99% sure that's a typo

Nnesha (nnesha):

right :=)

OpenStudy (anonymous):

it was a typo, thanks to both of you!

Nnesha (nnesha):

pleasure.

OpenStudy (usukidoll):

I think it's backwards though. with the quotient rule wouldn't it be \[\log_7(n)-\log_7(5)\]

OpenStudy (anonymous):

are you sure? 5 was the denominator

Nnesha (nnesha):

ohh yeah i thought n is the denominator

OpenStudy (usukidoll):

-_- I saw log7(5)-log7(n) that's 5/n not n/5

Nnesha (nnesha):

nvm arnv my bad i typed that wrong in the properties

OpenStudy (anonymous):

oh, don't worry about it, thank you guys again!

Nnesha (nnesha):

thanks! i'll write it correctly

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha quotient rule\[\large\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\] \(\color{blue}{\text{End of Quote}}\) correction quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

looks right ;)